OpenStudy (anonymous):

Help asap please? A beam of light passes through the air and enters a diamond at an angle of incidence of 40 degrees. Use Snell’s Law to find the angle of refraction of the diamond. (Index of refraction for air is 1.00 and the index of refraction of a diamond is 2.42.) If the beam of light passed through water at the same angle of incidence, would the angle of refraction from the diamond be the same and if not, by how much would it change? (Index of refraction for water is 1.33.) Show all work.

OpenStudy (anonymous):

Here's Snell's Law:\[n _{i}\sin \left( \theta _{i} \right)=n _{r}\sin \left( \theta _{r} \right)\]where ni is the index of refraction of the incident medium (air in this case); θi is the angle of incidence; nr is the index of refraction for the refracting material (diamond in this case); and θr is the angle of refraction. So since you need to find θr you just need to rearrange Snell's Law a bit to get this:\[\theta _{r}=\sin ^{-1}\left( \frac{ n _{i}\sin \left( \theta _{i} \right) }{ n _{r} } \right)\] Does that help?